Follow the flow of solving like the example given above.You may use the Kb and Kf of water in the example.

1.calculate the freezing and boiling point of a solution prepared by dissolving 15.5 g of Al(NO3)3 in 200.0 g of water.(Molar mass of Al(NO3)3 is 212.996 g/mol).

2. A solution is prepared by dissolving 120 grams of NaCl in 450 grams of water.Find the freezing and boiling point of this solution.(Molar mass of NaCl is 58.44 g/ mol).



1
Expert's answer
2021-10-21T02:21:52-0400

"Al(NO \n3\n\u200b\t\n ) \n3\n\u200b\t\n \u21d2Al \n3^+\n +3NO \n3^-\n\u200b"

Now here the mole of Aluminium Nitrate dissociates in 4 moles ,hence the value of (i=4)

"Formula used \Rightarrow delta T_f=i\times K_F \times m"


"where K_f=" 1.86 degre celcius kg mol-1


m=molality




Now :


Moles of aluminium nitrate

"\\frac{15.5}{212.996}\n\n\u200b\t\n =0.0728 moles"


"m=\dfrac{0.0728}{200}\times1000=0.364"


Now: For Freezing point:


"delta T_f=i\times K_f \times m"


"=4\times 1.86\times0.364=2.70"°C




Now we know that


"delta T_f={T^0}_f-T_f"


"Putting {T}^0_f=0"


We get the value of

"T_f=-2.70" degree celcius and this is our answer